Solve and check the equation for X
2 = 3 + 9
----- ------- -----------
x+2 x+6 x^2+8x+12
I am confused on this probelm. They added way to many things in there for me to understand. Thank you for the help

Answer by checkley71(8403)

2/(X+2)=3/(X+6)+9/(X^2+8X+12)
2/(X+2)=3/(X+6)+9/(X+6)(X+2)
2/(X+2)=[3(X+2)+9]/(X+6)(X+2) CANCEL OUT THE (X+2) IN BOTH DENOMINATORS THUS
2=[3(X+2)+9]/X+6) NOW CROSS MULTIPLY
2(X+6)=3(X+2)+9
2X+12=3X+6+9
2X-3X=-12+15
-X=3
X=-3
PROOF
2/(-3+2)=3/(-3+6)+9/(9-24+12)
2/-1=3/3+9/-3
-2=1-3
-2=-2
##############################################
MERRY CHRISTMAS